Print

Print


Blanch-SR1 CPU Issues
=====================


Ticket Details

 1360302                  
 Blanch-SR1 CPU Issues    
 open                     
 Mon Jul 09 08:30:00 2018 
 Not set                  
 NO                       

Problem Description

Cpu on Blanch-SR1 is hitting almost 100%


Problem Effects

There may be packet loss or latencies to some HEAnet services


Last Update Message

Mon Jul 09 18:50:48 2018

> We are continuing to investigate and carry out mitigations, we are
> planning to fail over to another switch in Blanchardstown to assist in
> troubleshooting.
The failover has completed successfully and has improved matters considerably.
We are monitoring overnight.


Affected Clients

 AGI, AIT, ANCHEIM, BCFE, CAO, CARLOWC, CDETB, CIT, COA, CONUL,           
    Cork-ETB, CORKETB, CSN, DCU, DFEI, DIAS, DIT, DKIT, EDUCAMPUS, EPA,      
    ESBS, ESRI, ETBI, FPA, FULBRIGHT, GCT, GMIT, GSI, HEA, HEANET, HRB,      
    HSE, IADT, ICHEC, IEDR, INEX, IPA, IPS, IRAAL, IRCHSS, IRCSET, ITB,      
    ITC, ITS, ITT, ITTD, IUA, JISC, KILLESTER, KWETB, LCETB, LIT, LMETB,     
    LOETB, LYIT, MAC, MCIRL, MI, MIC, MIC-SPT, MLAB, MU, NCAD, NCI, NCRI,    
    NDLR, NDRC, NFETLHE, NUI, NUIG, OIR, PAC, PDST, QQI, QUALIFAX, QUB,      
    RCPI, RCSI, RIA, RIAM, SCHOOLS, SMB, STJOHNS, TCD, TEACHNET, Teagasc,    
    THEA, UCC, UCD, UL, WIT                                                  

Transaction History

 Date Created              Message                                                                   
 Mon Jul 09 14:24:39 2018  CPU on blanch SR-1 is hitting almost 100% since approximately 08:30       
                              this morning. We are investigating the cause of this and possible         
                              mitigations. In the meantime, there may be latency issues or packet       
                              loss while accessing HEAnet services including Filesender, Media,         
                              Moodles, and IdP serviers. We will have another update at 15:00.          
 Mon Jul 09 14:59:03 2018  We are continuing to investigate possible mitigations for the issues      
                              we're seeing. We will update again at 16:00                               
 Mon Jul 09 16:00:04 2018  We have taken some steps to mitigate the issues we are seeing and are     
                              continuing to monitor the situation. We will have another update          
                              before 17:30                                                              
 Mon Jul 09 17:30:07 2018  We are continuing to investigate and carry out mitigations, we are        
                              planning to fail over to another switch in Blanchardstown to assist in    
                              troubleshooting.                                                          
 Mon Jul 09 18:50:47 2018  > We are continuing to investigate and carry out mitigations, we are >    
                              planning to fail over to another switch in Blanchardstown to assist in    
                              > troubleshooting. The failover has completed successfully and has        
                              improved matters considerably. We are monitoring overnight.